# Carbon-based light emitting defects in different layers of hexagonal boron nitride

### v1

@@ IV.2 Phonon sideband
When examining the electron-phonon coupling of the optical transition, the relevant phonon replicas are high-energy modes corresponding to bond stretching, typically associated with substitutional C defects [40]. Fig. 4A shows the defect CBCB–c, whose phonon sideband is representative of most of the defects studied here. The optical excitation induces the transfer of one electron from a C atom to the other if the defect is CNCB, or from/to the C atom to the valence/conduction band if the defect is CXCX. Either way, the C–B and/or C–N bonds become weaker or stronger, resulting in a bond stretching: a high-energy localized longitudinal phonon.
However, there is one defect with a markedly different photoluminescence spectrum, CBCB–a, that is, the CB defects are opposite to each other and separated by a N atom from an intermediate layer. This defect features a low-energy phonon replica shifted just meV, see Fig. 4. In this defect, the C atoms are slightly displaced out of the plane (emulating a C–C repulsion) in the ground state, and the phonon associated with such distortion is responsible for the 62 meV phonon replica. In Fig. 4B, C, the phonon with the most relevant partial Huang-Rhys factor is shown for two different dimers (CBCB–a and CBCB–c). This behavior will be discussed in greater detail in the next section. To the best of our knowledge, there are a couple of reports of defects in hBN with a similar low-energy phonon replica, the so-called yellow emitters [69], but these emitters lack spin activity. Although we can discard CBCB–a as yellow emitters, it is likely that the yellow emitters trigger an out-of-plane distortion, explaining their low-energy phonon replicas. The other report of SPEs with a similar PL was obtained in a hBN flake after a H2 plasma treatment [70].
@@ IV.3 Weakly coupled spin pairs
Several reports of SPEs exhibit optically detected magnetic resonance (ODMR) with a very small zero-field splitting (ZFS) in the range of the MHz [20, 71, 72, 73]. Such a small ZFS can be associated with weakly coupled defects [41] or defects with a triplet ground state but with a very small ZFS. Indeed, most CXCX defects studied here are expected to have a small ZFS since both C atoms are separated by nm and the ZFS decays as (in the vacuum). Note that a small ZFS is not incompatible with a large exchange energy, separating the singlet and triplet states. There is a general model aiming to explain the already mentioned ODMR results [42]. It mostly fits with the CNCB–() defects studied here: donor-acceptor pairs, singlet ground state, C–C distance of nm. However, here, we show that weakly interacting triplet states could also be found in hBN multilayers. Particularly, the ODMR measured by Scholten et al [41] effectively behaves as spin 1/2 defects, as the authors demonstrated by Rabi oscillation measurements. Such behavior could occur in defects like the ones from our study, an actual triplet or singlet, but once the exchange energy becomes weaker, that is, when defects are further away. Going closer to the dissociation limit is not only computationally expensive but, more importantly, DFT is known for failing to describe such a limit.
@@ V Conclusions
However, if the two C atoms in different hBN layers are substituting N or B, they have a triplet ground state. It is separated by more than eV from the singlet state, making these defects spin-active above room temperature. Their ZPL can be larger than when the defects form donor-acceptor pairs, eV. Again, these defects present the typical high-energy phonon replicas in their photoluminescence spectra.
When two CB atoms are on top of each other but separated by a hBN interlayer, we calculated a markedly different phonon sideband in the photoluminescence spectrum, with well-defined phonon replicas at meV. They correspond to out-of-plane rearrangement upon excitations. We are unaware of any measurement of such a spectrum for spin-active defects, but the so-called (spinless) yellow emitters have similar phonon replicas and could also correspond to out-of-plane distortions.
@@ Acknowledgements.
- Gottscholl et al. [2021] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Room temperature coherent control of spin defects in hexagonal boron nitride, Science Advances 7, 10.1126/sciadv.abf3630 (2021).

### v2

@@ IV.2 Phonon sideband
When examining the electron-phonon coupling of the optical transition, the relevant phonon replicas are high-energy modes corresponding to bond stretching, typically associated with substitutional C defects [40]. Fig. 4A shows the defect CBCB–, whose phonon sideband is representative of most of the defects studied here, Figures S2-S4 in the supplementary material have the photoluminescence spectra for all remaining configurations. The optical excitation induces the transfer of one electron from a C atom to the other if the defect is CNCB, or from/to the C atom to the valence/conduction band if the defect is CXCX. Either way, the C–B and/or C–N bonds become weaker or stronger, resulting in a bond stretching: a high-energy localized longitudinal phonon.
However, there are two defects with a markedly different photoluminescence spectra, CBCB– and CBCB–. In the former, the CB defects are opposite to each other and separated by a N atom from an intermediate layer. This defect features a low-energy phonon replica shifted just meV, see Fig. 4A. In this defect, the C atoms are slightly displaced out of the plane (emulating a C–C repulsion) in the ground state, and the phonon associated with such distortion is responsible for the 62 meV phonon replica. A similar scenario occurs in CBCB–. Fig. 4B, C, shows the phonon with the most relevant partial Huang-Rhys factor , for two different dimers, CBCB– and CBCB–, while in the former the electronic excitation induces are rearrangement restricted to the defect plane, in the later the dominant phonon modes are out of plane.
To the best of our knowledge, there are a couple of reports of defects in hBN with a similar low-energy phonon replica, the so-called yellow emitters [72], but these emitters lack spin activity. Although we can discard CBCB– and CBCB– as yellow emitters, it is likely that the yellow emitters trigger an out-of-plane distortion, explaining their low-energy phonon replicas. The other report of SPEs with a similar PL was obtained in a hBN flake after a H2 plasma treatment [73].
@@ IV.3 Weakly coupled spin pairs
Several reports of SPEs in hBN exhibit optically detected magnetic resonance (ODMR) with very small zero-field splittings (ZFS) in the MHz range [20, 74, 75, 76]. Such small ZFS values can arise either from weakly coupled, physically distinct defects [41] or from individual defects possessing a triplet ground state with an intrinsically small ZFS. Our calculated CXCX defects are candidates for both scenarios: when the two carbon atoms are separated by 0.4-0.7 nm, the ZFS within their triplet state is expected to be small, primarily due to the decay of the magnetic dipole-dipole interaction. It is important to note that a small ZFS within the triplet manifold is not incompatible with a large exchange energy () that separates the singlet and triplet states, a situation relevant for our CXCX defects at shorter C-C distances (see Fig. 2). At larger separations, as the defects become spin degenerate.
At this point, it is informative to estimate the ZFS of the CXCX defects where the constituent carbon atoms are separated by distances in the range of 0.4-1.0 nm; the magnetic dipole-dipole interaction is expected to be a primary contributor. Using the point-dipole approximation (, with MHz nm3, see Section V from the supplementary material), we estimate the ZFS parameter to range from approximately 78 MHz (for nm) up to 1.22 GHz (for nm). For a specific configuration like CBCB– with an inter-carbon distance of 0.67 nm, this estimation yields a value of approximately 260 MHz. These values, particularly at the larger end of the separation range, are substantially smaller than the multi-GHz ZFS typically observed for compact spin-1 defects like V in hBN [77], and fall within a range that could be consistent with ODMR experiments reporting MHz-scale features.
A general model by Robertson et al. [42] proposes that some of the experimentally observed ODMR signals arise from pairs of weakly coupled defects. While our CNCB– defects share some characteristics with the systems considered by this model, such as C-C distances around nm, their established singlet ground state (due to charge transfer) makes them unlikely candidates to be these pairs themselves.
This distinction motivates exploring other defect types. Here, we propose that our CXCX homonuclear defect pairs, provide a compelling alternative origin for such ODMR signatures. As discussed, when these CXCX defects are sufficiently far apart, the exchange energy () diminishes, and the system behaves as two weakly interacting centers. The resulting state would exhibit a very small ZFS, as argued above. Such a system could indeed produce ODMR signals that appear to arise from ‘effectively defects,’ consistent with the experimental findings of Scholten et al. [41], who observed this characteristic behavior through Rabi oscillation measurements. Thus, the CXCX pairs at in hBN multilayers naturally reconcile the picture of weakly interacting spins with an underlying triplet state possessing a minimal ZFS.
IV.4 Suitability or usefulness as SPEs
Overall, the studied defects present a mix of typical behaviors for in-plane defects and new features. Heteronuclear CBCN defect pairs are predicted to behave as the typical bright SPEs usually found in hBN [1]. They have a typical phonon sideband with high-energy replicas (160-190 meV), and their emission covers the entire visible range, also they usually do not produce a measurable ODMR signal [28]. All these characteristics suggest a large family of related defects, as proposed by different arrangements of substitutions C defects acting as donor-acceptor systems [40, 25]. Our contribution shows that in such a family, C defects residing in different layers are also candidates for SPEs.
Instead, the homonuclear defects CBCB, CNCN show some interesting features for SPEs, (i) often they are spin-active at room temperature (i.e., a triplet) and (ii) regardless of details, each series (CBCB or CNCN) has a very similar ZPL, converging to an asymptotic limit. Both features are not present in C dimers within the same monolayer [31] and are interesting for SPEs. On one hand, most of the already known spin-active defects in hBN are based on vacancies [35], which share many favorable and detrimental aspects with well-studied SPEs such as the NV-center in diamond (e.g., triplet ground state, ZFS in the GHz range, broad emission line). Instead, most defects spin-active CBCB, CNCN defects would have some interesting properties as SPEs: a bright emission, a narrow emission line, and well-separated phonon replicas. But, their ZFS would be non-constant and range from tens to hundreds of MHz, which could make the integration with current technology difficult.
There are two particular defect configurations, CBCB– and CBCB–, with a phonon sideband different, to the best of our knowledge, from other substitutional C-based SPEs, characterized by prominent meV replicas from out-of-plane modes. If this distinct vibronic structure leads to an effectively broader overall emission profile where the ZPL is less dominant or more difficult to spectrally isolate, it could hinder applications requiring high photon indistinguishability and long coherence times, as these rely heavily on the properties of the ZPL photons.
Arguably, the most significant problem of SPEs in hBN is the difficulty of making monochromatic SPEs within the visible. So far, there are two irradiation protocols to achieve monodisperse emitters: the so-called blue [78] and the yellow [72] emitters. In this respect, very specific defect geometries -such as CBCB–- are unlikely to be deterministically created. Instead, there are irradiation protocols aimed at producing defects [79]. Such protocols could be adapted [80] to produce pairs of within less than a nm. In the presence of carbon contamination [4], these defects could turn into CBCB dimers as those studied here, some of them with a triplet ground state.
Are the pairs of C defects considered here SPEs at all? For strongly coupled CXCX pairs forming a triplet state, or for CNCB singlet pairs, the pair acts as a single quantum emitter. In the limit of weakly interacting CXCX pairs, while the two centers are nearly independent, single-photon emission would still be expected if the optical cycle involves defined joint states of the pair. Definitive confirmation of single-photon emission characteristics relies on measurements. As stated before, measurements of hBN defects of yet undetermined atomic structure reveal a very small ZFS tens of MHz or even smaller. These reports routinely show the dip associated with single-photon emission. This is a strong hint that well-separated defects can behave as SPEs. Particularly, one article [20] shows a defect in hBN with a ZFS of 50 MHz, together with a photoluminescence spectrum featuring the typical high-energy phonon replicas (180 meV) of C-based defects in hBN and a ZPL of eV. All these measurement points to CNCN defects of the type studied here, perhaps with a larger C–C distance (1.1-1.2 nm).
Even though we didn’t calculate the emission rates of the defects studied here, it is relevant to ask if these weakly interacting pairs have an emission rate relevant for applications (or even detection). We can not give a definitive answer. However, the experiments mentioned above suggest a large enough emission rate, as the single defects provide a clear signal. For the donor-acceptor-like pairs, the charge transfer should imply a large dipole transition moment and a large emission rate. Experimentally, their “brightness is the highest reported for a quantum emitter in the visible spectral range” [1].
@@ V Conclusions
However, if the two C atoms in different hBN layers are substituting the same species, N or B, they likely have a triplet ground state for closer distances Å. It is separated by more than eV from the singlet state, making these defects spin-active above room temperature. Their ZPL is almost independent of the C-C distance (i.e. nearly monochromatic), according to SCAN-based calculations, their emission is within the red. Calculations with the HSE06 functional shift the emission energy to the yellow-to-green range. Regardless of the actual value of the ZPL, the nearly monochromatic emission holds. Most of these kinds of defect dimers present the typical high-energy phonon replicas in their photoluminescence spectra.
As the distance of the mononuclear defects (CBCB or CNCN) increases, they become practically spin degenerate, or ‘weakly interacting spin pairs’ [42]. In either case, a robust triplet or a spin degenerate state, the defects are well-separated and their zero field splitting should be within the MHz range, which is consistent with some optically detected magnetic resonance (ODMR) experiments [41, 20, 74, 75, 76].
When two CB atoms are on top of each other but separated by a hBN interlayer or lying as close as possible in adjacent layers -CBCB– and CBCB-- their photoluminescence spectrum has a markedly different phonon sideband, with well-defined phonon replicas at meV. They correspond to out-of-plane rearrangement upon excitations. We are unaware of any measurement of such a spectrum for spin-active defects, but the so-called (spinless) yellow emitters [72] have similar phonon replicas and could also correspond to out-of-plane distortions.
@@ Acknowledgements.
- Gottscholl et al. [2021a] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Room temperature coherent control of spin defects in hexagonal boron nitride, Science Advances 7, 10.1126/sciadv.abf3630 (2021a).
@@
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
@@
- Perdew [1985] J. P. Perdew, Density functional theory and the band gap problem, International Journal of Quantum Chemistry 28, 497 (1985).
- Reimers et al. [2018] J. R. Reimers, A. Sajid, R. Kobayashi, and M. J. Ford, Understanding and calibrating density-functional-theory calculations describing the energy and spectroscopy of defect sites in hexagonal boron nitride, Journal of Chemical Theory and Computation 14, 1602 (2018).
@@
- Gottscholl et al. [2021b] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, D. Krauße, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Spin defects in hbn as promising temperature, pressure and magnetic field quantum sensors, Nature Communications 12, 4480 (2021b).
- Gale et al. [0] A. Gale, C. Li, Y. Chen, K. Watanabe, T. Taniguchi, I. Aharonovich, and M. Toth, Site-specific fabrication of blue quantum emitters in hexagonal boron nitride, ACS Photonics 0, null (0).
- Carbone et al. [2025] A. Carbone, I. D. Breev, J. Figueiredo, S. Kretschmer, L. Geilen, A. B. Mhenni, J. Arceri, A. V. Krasheninnikov, M. Wubs, A. W. Holleitner, A. Huck, C. Kastl, and N. Stenger, Quantifying the creation of negatively charged boron vacancies in he-ion irradiated hexagonal boron nitride (2025), arXiv:2501.18481 [cond-mat.mes-hall] .
- Ren et al. [2025] F. Ren, Z. Xu, and Y. Wu, Optimization of carbon irradiation parameters for creating spin defects in hexagonal boron nitride, Nanotechnology and Precision Engineering 8, 033014 (2025), https://pubs.aip.org/tu/npe/article-pdf/doi/10.1063/10.0036119/20472607/033014_1_10.0036119.pdf .